### v2

@@ Appendix D Proof of Section 5
Next, we show the optimal allocation of the inf-convolution exists if and only if the minimizer of (9) exists. Suppose $(X_{1},X-X_{1})$ is the optimal allocation of the inf-convolution. Then we have $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\square\rho^{\mathbb{Q}_{2}}(X)$. Let $x_{1}=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})$ and $y_{1}>x_{1}\vee\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$. We choose $U_{X_{1}}^{\mathbb{Q}_{1}}$ such that $X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ for all $w\in\Omega$. Using the above argument, we have $F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})\leqslant\Lambda_{x_{1},y_{1}}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$. By the monotonicity of $\rho$, we have $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})\geqslant x_{1}+\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x_{1},y_{1}}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}}))$. This implies $x_{1}+\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x_{1},y_{1}}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}}))=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}$. Hence, $(x_{1},y_{1},U_{X_{1}}^{\mathbb{Q}_{1}})$ is the minimizer of (9). Moreover, if $(x_{1},y_{1},U)$ is the minimizer of (9), one can easily check that $(\Lambda_{x_{1},y_{1}}^{-1}(U),X-\Lambda_{x_{1},y_{1}}^{-1}(U))$ is the optimal allocation of the inf-convolution. We complete the proof. ∎

### v3

@@ Appendix A Proof of all results in Sections 2-5
Next, we show for $\mathcal{X}=L^{\infty}$, the optimal allocation of the inf-convolution exists if and only if the minimizer of (10) exists. Suppose $(X_{1},X-X_{1})$ is the optimal allocation of the inf-convolution. Then we have $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\square\rho^{\mathbb{Q}_{2}}(X)$. Let $x_{1}=\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})$ and $y_{1}>x_{1}\vee\mathrm{ess\mbox{-}sup}^{\mathbb{Q}_{1}}X_{1}$. We choose $U_{X_{1}}^{\mathbb{Q}_{1}}$ such that $X_{1}\leqslant F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$ for all $w\in\Omega$. Using the above argument, we have $F_{X_{1}}^{\mathbb{Q}_{1},-1}(U_{X_{1}}^{\mathbb{Q}_{1}})\leqslant\Lambda_{x_{1},y_{1}}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}})$. By the monotonicity of $\rho$, we have $\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}(X_{1})+\rho^{\mathbb{Q}_{2}}(X-X_{1})\geqslant x_{1}+\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x_{1},y_{1}}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}}))$. This implies $x_{1}+\rho^{\mathbb{Q}_{2}}(X-\Lambda_{x_{1},y_{1}}^{-1}(U_{X_{1}}^{\mathbb{Q}_{1}}))=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}_{1}}{\sim}U[0,1]}\left\{x+\rho^{\mathbb{Q}_{2}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}$. Hence, $(x_{1},y_{1},U_{X_{1}}^{\mathbb{Q}_{1}})$ is the minimizer of (10). Moreover, if $(x_{1},y_{1},U)$ is the minimizer of (10), one can easily check that $(\Lambda_{x_{1},y_{1}}^{-1}(U),X-\Lambda_{x_{1},y_{1}}^{-1}(U))$ is the optimal allocation of the inf-convolution. We complete the proof. ∎
Proof of Corollary 8. In light of Theorem 6, we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}}\mathop{\square}\displaylimits\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}(X)=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x}\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\left\{x+\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}\left(X-\Lambda_{x,y}^{-1}(U)\right)\right\}.$
Let $\hat{F}_{x,y}(z)=\{F^{\mathbb{Q}}_{X-\Lambda_{x,y}^{-1}(U)}(z):U\overset{\mathbb{Q}}{\sim}U[0,1]\}$ for $z\in\mathbb{R}$ and $\overline{\mathrm{VaR}}_{\beta}^{\mathbb{Q}}$ represent $\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}$ for $\Lambda_{1}=\beta\in(0,1)$. This means $\overline{\mathrm{VaR}}_{\beta}^{\mathbb{Q}}$ is the right $\mathrm{VaR}$. By Theorem 1 of Han and Liu (2025), we have
$\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\Lambda_{1}\mathrm{VaR}}^{\mathbb{Q}}\left(X-\Lambda_{x,y}^{-1}(U)\right)=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y}).$
It follows from Corollary 1 of Han and Liu (2025) that for $\alpha\in(0,1)$,
$q_{\alpha}^{+}(\hat{F}_{x,y})=\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{\alpha}^{\mathbb{Q}}(X-\Lambda_{x,y}^{-1}(U)),$
where $q_{\alpha}^{+}$ is $q_{\Lambda_{1}}^{+}$ with $\Lambda_{1}=\alpha$. Direct computation shows
$\inf_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{\alpha}^{\mathbb{Q}}(X-\Lambda_{x,y}^{-1}(U))=-\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\mathrm{VaR}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U)).$
In light of Lemma 4.3 of Bernard et al. (2014) and the result in Makarov (1981), we have
$\sup_{Y\overset{\mathbb{Q}}{\sim}F_{X}^{\mathbb{Q}},U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-Y+\Lambda_{x,y}^{-1}(U))$ $=\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(V))$
$\leqslant\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U)),$
where $V=U_{-X}^{\mathbb{Q}}\mathds{1}_{\{U_{-X}^{\mathbb{Q}}\leqslant\alpha\}}+(1+\alpha-U_{-X}^{\mathbb{Q}})\mathds{1}_{\{U_{-X}^{\mathbb{Q}}>\alpha\}}$, which implies
$\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U))$ $=\sup_{Y\overset{\mathbb{Q}}{\sim}F_{X}^{\mathbb{Q}},U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-Y+\Lambda_{x,y}^{-1}(U))$
$=\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(V)).$
Using the above conclusion and the continuity of $\Lambda_{x,y}^{-1}$ and $F_{X}^{\mathbb{Q},-1}$, applying Lemma 4.4 of Bernard et al. (2014), we have
$\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\mathrm{VaR}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U))=\sup_{U\overset{\mathbb{Q}}{\sim}U[0,1]}\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(U)).$
Hence,
$q_{\alpha}^{+}(\hat{F}_{x,y})=-\overline{\mathrm{VaR}}_{1-\alpha}^{\mathbb{Q}}(-X+\Lambda_{x,y}^{-1}(V))=l_{x,y}(1-\alpha),$
where $l_{x,y}(t)=\sup_{s\in(0,t]}\{F_{X}^{\mathbb{Q},-1}(s)-\Lambda_{x,y}^{-1}(1-t+s)\},~t\in(0,1)$. Direct computation shows $\hat{F}_{x,y}(z)=l_{x,y}^{-1}(z)$ on $\mathbb{R}$ except a countable number of points.
Next, we show $q_{\Lambda_{1}}^{+}(l_{x,y}^{-1})=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y})$. Let $z_{0}=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y})$. Clearly, $z_{0}\in\mathbb{R}$. By definition, there exist $z_{n}\downarrow z_{0}$ such that $\hat{F}_{x,y}(z_{n})>1-\Lambda_{1}(z_{n})$. It follows from the right continuity of $\Lambda_{1}$ and the monotonicity of $\hat{F}_{x,y}$ that there exist $0<\varepsilon_{n}<1/n$ such that $\hat{F}_{x,y}(z)>1-\Lambda_{1}(z)$ for $z\in(z_{n},z_{n}+\varepsilon_{n})$. Then there exist $w_{n}\in(z_{n},z_{n}+\varepsilon_{n})$ such that $l_{x,y}^{-1}(w_{n})=\hat{F}_{x,y}(w_{n})>1-\Lambda_{1}(w_{n})$. Hence, $q_{\Lambda_{1}}^{+}(l_{x,y}^{-1})\leqslant w_{n}\to z_{0}=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y})$. We can similarly show the inverse inequality. Hence, $q_{\Lambda_{1}}^{+}(l_{x,y}^{-1})=q_{\Lambda_{1}}^{+}(\hat{F}_{x,y})$. We complete the proof. ∎
@@
Proof of Corollary 9. In light of Proposition 6, we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)$ $=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x\vee x_{1}}\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\left\{x+\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X-x\mathds{1}_{B_{1}}-x\vee x_{1}\mathds{1}_{B_{2}}-y\mathds{1}_{(B_{1}\cup B_{2})^{c}}\right)\right\}$
$=\inf_{x\in\mathbb{R}}\inf_{y\geqslant x\vee x_{1}}\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X+0\wedge(x-x_{1})\mathds{1}_{B_{2}}+(x-y)\mathds{1}_{(B_{1}\cup B_{2})^{c}}\right)$
$=\inf_{(B_{1},B_{2})\in\mathcal{B}_{\lambda_{1},\lambda_{2}}}\lim_{y\to\infty}\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X-y\mathds{1}_{B_{1}^{c}}\right).$
From the above expression, we should choose $B_{1}$ such that $\mathbb{Q}_{2}(B_{1})$ is as small as possible. Hence, if $\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})\geqslant\alpha\mathbb{P}(D_{2})$, the optimal way to choose $B_{1}$ is to set $B_{1}=(D_{1}\cap D_{2}^{c})\cup(\{U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2})$ with $\mathbb{P}(\{U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2})=\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})$. Then we have
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=\lim_{y\to\infty}\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X-y\mathds{1}_{B_{1}^{c}}\right)=\frac{1}{\alpha}\mathbb{E}^{\mathbb{Q}_{2}}\left(X\mathds{1}_{\{t^{-}<U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2}}\right)$
with $\mathbb{P}(\{t^{-}<U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2})=\alpha\mathbb{P}(D_{2})$. Note that $\frac{d\mathbb{Q}_{2}}{d\mathbb{Q}}=\frac{\mathbb{P}(D_{1}\cap D_{2})}{\mathbb{P}(D_{2})}$ on $D_{1}\cap D_{2}$. Hence, we have
$\mathbb{E}^{\mathbb{Q}_{2}}\left(X\mathds{1}_{\{t^{-}<U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2}}\right)$ $=\frac{\mathbb{P}(D_{1}\cap D_{2})}{\mathbb{P}(D_{2})}\mathbb{E}^{\mathbb{Q}}\left(X\mathds{1}_{\{t^{-}<U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2}}\right)$
$=\frac{\mathbb{P}(D_{1}\cap D_{2})}{\mathbb{P}(D_{2})}\int_{\alpha_{0}}^{\alpha_{1}}\mathrm{VaR}_{t}^{\mathbb{Q}}(X)\mathrm{d}t.$
with $\alpha_{0}=1-\frac{\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})}{\mathbb{P}(D_{1}\cap D_{2})}$ and $\alpha_{1}=1-\frac{\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})-\alpha\mathbb{P}(D_{2})}{\mathbb{P}(D_{1}\cap D_{2})}$, which implies the desired result.
If $\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})\leqslant 0$, we choose $B_{1}\subset D_{1}\cap D_{2}^{c}$. Then $\mathbb{Q}_{2}(B_{1}^{c})=1$, which implies $\Lambda\mathrm{VaR}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=$-$\infty$.
If $0<\lambda_{1}\mathbb{P}(D_{1})-\mathbb{P}(D_{1}\cap D_{2}^{c})<\alpha\mathbb{P}(D_{2})$, then we choose $B_{1}=(D_{1}\cap D_{2}^{c})\cup(\{U_{X}^{\mathbb{P}}<t^{+}\}\cap D_{1}\cap D_{2})$, where $t^{+}$ is defined above. Direct computation shows $\mathbb{Q}_{2}(B_{1})<\alpha$, which implies
$\Lambda\mathrm{VaR}^{+,\mathbb{Q}_{1}}\mathop{\square}\displaylimits\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}(X)=\lim_{y\to\infty}\mathrm{ES}_{\alpha}^{\mathbb{Q}_{2}}\left(X-y\mathds{1}_{B_{1}^{c}}\right)=-\infty.$
This completes the proof. ∎